OpenStudy (zenmo):

For what t-values is the curve concave upward? x = 2sint, y = 3cost, 0 < t < 2pi

OpenStudy (zenmo):

\[\frac{ dy }{ dx }=\frac{ -3 }{ 2 }\tan(t)\]\[\frac{ d^2y }{ dx^2 }=\frac{ -3 }{ 4 }\sec^3(t)\]\[\frac{ -3 }{ 4 }\sec^3(t) > 0\]\[\sec^3(t)<0\] I am stuck

OpenStudy (zenmo):

All of the steps are correct, so you don't have to verify the work.

OpenStudy (amorfide):

I would find t values for which dy/dx=0 then I would inspect it below and above those values, to see if it is concave up this is only a suggestion since there may be a more efficient way

OpenStudy (amorfide):

but peresonally x = 2sint, y = 3cost x/2=sint y/3=cost square both (x/2)^2=sin^2t (y/3)^2=cos^2t add together sin^2t+cos^2t=1 (x/2)^2 + (y/3)^2=1 maybe this can help somehow, I gotta go to sleep good night!

zepdrix (zepdrix):

Forgot which quadrants your secant is negative in? Can convert to cosines if that helps. \(\large\rm \frac{1}{\cos^3t}<0\) Cosine is our x-coordinate as we're traveling along the unit circle, ya? So I guess we're negative in quadrants 2 and 3.

OpenStudy (zenmo):

so the interval would be ( pi/2, 3pi/2 )

zepdrix (zepdrix):

Mmmmmmmmm ya that sounds right.

OpenStudy (zenmo):

would I state the interval as concave up or down?

OpenStudy (zenmo):

Since the equality sign got changed, and the question is asking for concave up.

OpenStudy (zenmo):

or do I just leave it as (pi/2, 3pi/2) and not state the concavity

zepdrix (zepdrix):

\[\large\rm y''=-\frac34\sec^3t\]And you made the statement that this expression needs to be larger than 0 for `concave up`.\[\large\rm 0<-\frac34\sec^3t\] The only way we can get concave up is if sec^3t is negative, (That makes the entire thing positive, right?)

OpenStudy (zenmo):

yes

zepdrix (zepdrix):

So you found the interval for `concave up`. Don't worry about the inequality flip and all that fancy stuff that happened in between. It didn't change what we were looking for.

OpenStudy (zenmo):

Ah I see, thanks for the rescue :)

zepdrix (zepdrix):

yay team!